# Heterogeneous Treatment Effects and

### v3

@@ Appendix G Strengthening Assumptions
Figure A.5: DGP for an indirectly-affected outcome, $h(Y)$.
G.1 Discrete Outcomes under Monotonicity Assumptions
@@
In practice, people frequently and implicitly assume monotonicity of treatment effects. We ask whether this assumption permits inference about mechanism activation for indirectly-affected outcomes of interest. To be specific, consider the two DGPs in figure A.6. We will index these DGPs by $s\in\{1,2\}$ where $s=1$ corresponds to the left DAG and $s=2$ corresponds to the right DAG.
$Z$$M$$Y_{1}$$h(Y_{1})$$X$$Z$$M$$Y_{2}$$h(Y_{2})$$X$
Figure A.6: Two different DGPs. On the left, in DGP 1, $X$ is a MDV. On the right, in DGP 2, $X$ is not a MDV.
@@
The left panel assumes $X_{k}$ is a MDV. $X_{k+1}$ is not a MDV in the right panel. In the figure, there are no other mediators. Therefore, both graphs satisfy exclusion assumptions 1 and 2 by construction.
@@
$h(Y_{s})$ =$=\begin{cases}0&Y_{s}\in(-\infty,c]\\ 1&Y_{s}\in(c,\infty],\end{cases}$ (56)
for some $c\in(-\infty,\infty)$. Propositions 3 and 4 show that we cannot differentiate between the DGPs 1 and 2 in Figure A.6 on the basis of the existence or non-existence of HTEs for indirectly-affected outcome $h(Y_{s})$.
We will consider what can be gained by imposing a monotonicity assumption of the form: $\frac{\partial^{2}Y}{\partial Z\partial X}>(<)0$ (note that the inequalities are strict). [^20] Writing the monotonicity assumption in this way assumes that this derivative exists. Clearly montonicity can hold in the left panel (where $X$ is a MDV) but $\frac{\partial^{2}Y}{\partial Z\partial X}=0$ in the right panel in which $X$ is not a MDV. To explore the implications of monotonicity, we consider following DGPs for $s=\{1,2\}$:
$Y_{i}=g_{i}(Z,X)+e$ (57)
@@
We ask whether researchers can differentiate these two cases when motonicity holds for the first DGP (e.g., an assumption of monotonicity). First, given equation (56) and equation (57), note that:
$E[h(Y_{s})|Z,X]=\Pr(e\geq c-g_{s}(Z,X))$ (58)
@@
$-f^{\prime}_{e}(c-g_{2})\frac{\partial g_{2}}{\partial X}\frac{\partial g_{2}}{\partial Z}$ (60)
The additional term $-f_{e}(c-g_{1})\beta(x,z)$ in equation (59) may help us to differentiate two DGPs by generating a differently-signed HTE.
@@
If, under certain $x$ and $z$, equation (60) and the first term of equation (59) have the same sign and the second term of equation(59) has the opposite sign and is sufficiently large, equation (60) and equation (59) will have different signs. Moreover, If $e$ is uniformly distributed, then $f^{\prime}_{e}=0$ and thus equation (60) is equal to 0 while equation (59) is non-zero.
@@
Consider the indirectly-affected outcome $h(Y_{s})$ satisfying equation (56) in which moderation effect $\beta(x,z)$ is monotonic.
@@
Proposition 10 reveals and Corollary A1 illustrates that Assumptions 2-3 and monotonicity are not, in general, jointly sufficient to use HTEs to assess mechanism activation on indirectly-affected outcomes.
G.2 Assumptions on the Latent Utility Distributions
In practice, one of the most common cases is that $Y$ is the latent utility, and $h(Y)$ is the observed action or discrete choice. If we can re-construct the utility from the observed data, then we can use it as the directly-affected outcome and use HTEs to assess mechanism activation under Assumptions 1-2. From Propositions 1 and 2, we know more information on mechanism activation can be ascertained from HTEs on directly-affected outcome.
@@ G.2 Assumptions on the Latent Utility Distributions
In practice, utility is usually specified as a linear function $V^{i}_{m}=X\beta$, where $X$ is the vector of observed variables and $\beta$, are parameters. We can estimate $\beta$ from the data. Then, we could treat $V_{m}^{i}$ as the directly-affected outcome to explore the mechanism using HTE. It should be clear that Propositions 1 and 2 hold for $V_{i}^{m}$ even if the utility is $U$.
Now, we use our motivating example to illustrate how to apply the above RUM and distribution assumptions in practice. In the main text, we provide a simple model of the effect of exogenous shocks on pro-incumbent voting. The model specifies the systematic component of utility, $V_{m}^{i}$. A RUM requires us to specify a random component of utility, $\varepsilon$ which also affects individual behavior and is additively separable from $V_{m}^{i}$ (i.e. not relevant to the theory of interest). Therefore, given the observed voting data we have, if we assume the distribution of $\varepsilon$ is type-I extreme value, then the probability individual votes for the incumbent is
@@
In practice, researchers tend to specify a linear model to approximate $V_{m}^{i}$, though other functional forms are also possible. If $V_{m}^{i}$ is correctly specified, we then treat $V_{m}^{i}$ as a directly-affected outcome (utility) and use this measure when estimating HTE.
G.3 Assumptions about the Magnitude of HTEs
@@
$\mathbb{P}[M_{1}\text{active}\mid\text{as extreme as HTE}]$ $=\frac{p\mathbb{P}[\text{HTE}|\text{active}]}{p\mathbb{P}[\text{HTE}|\text{active}]+(1-p)\mathbb{P}[\text{HTE}|\text{not active}]}$
@@
Figure A.8: Illustration of prior and posterior beliefs under a model in which $HTE=5$,$\mu_{1}=1$, $\mu_{2}=0.5$, $Beta(2,3)$, and $\sigma^{2}_{1}=\sigma^{2}_{2}=1$. The densities reflect the prior and posterior distributions of $p$, the probability that mechanism $1$ is active. The shift in the distribution to the right from prior to posterior suggests that the large HTE strengthens our belief that mechanism $1$ is active.
@@
Appendix H Simulation
Illustration: The distinction between directly-affected and indirectly-affected outcomes is novel to this paper. To illustrate the logic and implications of learning about mechanisms from HTEs in the case of an indirectly-affected outcome, we provide a short simulation that incorporates real attitudinal data. Specifically, we consider a hypothetical persuasion experiment that aims to shift support for greenhouse gas regulation among partisans in the US. Consistent with approaches used by scholars of persuasion, we will examine heterogeneity in partisan affiliation (here, simplified to Democrats and Republicans) (see Coppock, 2022: etc.). We use data on (1) partisan affiliation; (2) support for greenhouse gas regulation, coded as a binary outcome where $1$ designates support for increased regulation; and (3) demographic covariates from the 2020 American National Election Study. It is useful to note that partisans’ opinions are relatively polarized on this issue: while 82.2% (95% CI: [80.5%, 84.0%]) of Democrats favor increasing regulations, just 38.3% (95% CI: [35.7%, 40.9%]) of Republicans favor such regulations. This suggests that partisanship is strongly prognostic of support for greenhouse gas regulation.
Various theories of learning or additudinal change incorporate mechanisms that imply that partisans may react differently to information about greenhouse gas regulation. Little, Schnakenberg, and Turner (2022) classify two mechanisms for belief formation and attitude change: accuracy and directional motives. Within their model, ideology (partisanship) is posited as a moderator of directional motives but not accuracy motives, meaning that partisanship is a candidate MDV for directional motives. To this end, we simulate different processes of attitudinal change to examine when we observe HTEs in partisanship. Our simulation proceeds as follows:
[…]
We vary $\tau\in[-1.5,1.5]$, which are treatment effects on a logistic scale. [^21] The assumption of a constant $\tau$ is clearly a simplification for the purposes of illustration. It does not generally follow from the Little, Schnakenberg, and Turner (2022) model. Figure A.9 reports the results of our simulation. In the left panel, we see that for non-zero treatment effects (e.g., for any $\tau\neq 0$), we always observe HTEs in partisanship, even when effects on the latent scale are homogeneous, e.g., the degree of attitudinal change is not moderated by partisanship. We observe different treatment effects for Democrats and Republicans on the binary outcome even with homogeneous treatment effects on the latent attitude because of different densities of respondents about the relevant cutpoint in the latent variable (see Figure A.10).
Figure A.9: Note that $N=2883$ partisans. We assess power at the $\alpha=0.05$ level.
Figure A.10: Distribution of latent $Y_{i}(0)$’s, by party.

### v4

@@ G.2 Justification of theoretical representations
    This framework most closely aligns with the model we present as a running example, though it omits a valence shock.
  2. 2.
    Arias et al. (2019) present a formal model in which the voter’s utility is given by their equations (1) and (3):
    $u_{i}(I\mid s)$ $=\underbrace{-(\pi^{1}(s)\theta^{H}+[1-\pi^{1}(s)]\theta^{L})}_{\text{Expressive disutility from party }p}+\underbrace{\delta_{i}}_{\text{Partisan bias}}+\underbrace{I[s\in\{l,h\}]I[\Delta e(s)\leq 0]\sum_{j\in N_{i}}I[\Delta u_{j}(I\mid s)\geq 0]}_{\text{Coordination motive}}$
    Here, the signal, $s$, is corruption information ($c_{i}$ in our notation). Given the assumption of common prior that the incumbent is a high type ($\theta^{H})$, the expressive disutility on the basis of the posterior $\pi^{1}(s)$ is analogous to our distaste for incumbent corruption $\lambda_{i}c_{i}$. They include a second mechanism, the voter’s coordination motives, which depends on the signal ($s$). This is absent from our motivating model. Partisan bias, $\delta_{i}$, is assumed to be a uniformly-distributed random variable. In our model, this is analogous to $a_{i}$. Importantly, partisan bias is not assumed to moderate either mechanism in our model or in theirs. Further, the coordination and distaste mechanisms are assumed to be additively separable with respect to utility.
  3. 3.
    Anduiza, Gallego, and Muñoz (2013) present a verbal argument about the possible channels through which partisanship might influence assessments of a candidate’s corruption. They argue that:
    “The limited information hypothesis (or the ignorant voter in its earliest version) argues that citizens vote for corrupt politicians because they are not sufficiently informed about their misbehaviors. If people knew with certitude that an official was corrupt, they would evaluate that person negatively and vote against him or her.” (p. 1677).
    This suggests that provision of information about corruption ($c_{i}$) should link to a citizens’ assessment of incumbent corruption. But, they note that:
    “The literature suggests two alternative explanations: citizens do not give credibility to this information or they do not consider it important enough compared with other elements when they think about politics.” (p. 1677)
    The importance qualification for the effect of information provision connects to corruption aversion (i.e. $\lambda_{i}$)—how much does corruption matter to the voter? The credibility qualification is subsequently associated with motivated reasoning:
[…]
    Motivated information processing/reasoning is subsequently connected to partisan affiliation, $a_{i}$ as a moderator of the effect of information on assessments of the incumbent’s corruption.
  4. 4.
    de Figueiredo, Hidalgo, and Kasahara (2023) present a verbal argument that first emphasizes the effect of information on voting behavior:
    “An important precondition for electoral accountability is whether or not voters have access to information about the corrupt behavior of public officials, which may prompt them to vote against such candidates on Election Day.” (p. 729)
    This justifies the link between information disclosure ($c_{i}$) and assessments of incumbent corruption. They subsequently argue that partisan bias can condition the effects of information on voting behavior:
    “Partisan bias may function as a buffer against corruption scandals and hinder electoral accountability for voters in democracies with established parties.” (p. 729)
    They argue that partisan bias shapes corruption aversion, which in turn leads to variation in response to corruption information and subsequent voter behavior.
    “Relatively young parties that built their reputation among voters on an anti-corruption platform may thus suffer harsher punishment from sympathizers when compared to other traditional parties.” (p. 730)
Appendix H Strengthening Assumptions
In the main text, Propositions 4 and 5 indicate that the existence or non-existence of HTEs are not generally informative about mechanism activation. In this section, we explore the conditions under which invoking stronger assumptions can provide more information.
Recall the basic problem with transformed outcomes. In Figure A.8, $h(Y)$ is the non-linear transformation of $Y$. Our main result shows that $X_{k+1}$ can also induce HTEs even though $X_{k+1}$ does not detect the mechanism.
[…]
Figure A.8: DGP for a transformed outcome, $h(Y)$.
H.1 Transformed Outcomes under Monotonicity Assumptions
@@
In practice, people frequently and implicitly assume monotonicity of treatment effects. We ask whether this assumption permits inference about mechanism activation for transformed outcomes of interest. To be specific, consider the two DGPs in figure A.9. We will index these DGPs by $s\in\{1,2\}$ where $s=1$ corresponds to the left DAG and $s=2$ corresponds to the right DAG.
[…]
Figure A.9: Two different DGPs. On the left, in DGP 1, $X$ is an MDV. On the right, in DGP 2, $X$ is not an MDV.
@@
The left panel assumes $X_{k}$ is an MDV. $X_{k+1}$ is not an MDV in the right panel. In the figure, there are no other mediators. Therefore, both graphs satisfy exclusion Assumptions 1 and 2 by construction.
@@
$h(Y_{s})$ $=\begin{cases}0&Y_{s}\in(-\infty,c]\\ 1&Y_{s}\in(c,\infty],\end{cases}$ (47)
for some $c\in(-\infty,\infty)$. Propositions 4 and 5 show that we cannot differentiate between the DGPs 1 and 2 in Figure A.9 on the basis of the existence or non-existence of HTEs for outcome $h(Y_{s})$.
We will consider what can be gained by imposing a monotonicity assumption of the form: $\frac{\partial^{2}Y}{\partial Z\partial X}>(<)0$ (note that the inequalities are strict). [^22] Writing the monotonicity assumption in this way assumes that this derivative exists. Clearly montonicity can hold in the left panel (where $X$ is an MDV) but $\frac{\partial^{2}Y}{\partial Z\partial X}=0$ in the right panel in which $X$ is not an MDV. To explore the implications of monotonicity, we consider following DGPs for $s=\{1,2\}$:
$Y_{i}=g_{i}(Z,X)+e_{i}$ (48)
@@
We ask whether researchers can differentiate these two cases when motonicity holds for the first DGP (e.g., an assumption of monotonicity). First, given (47) and (48), note that:
$E[h(Y_{s})|Z,X]=\Pr(e_{i}\geq c-g_{s}(Z,X))$ (49)
@@
$f^{\prime}_{e}(c-g_{2})\frac{\partial g_{2}}{\partial X}\frac{\partial g_{2}}{\partial Z}$ (51)
The additional term $f_{e}(c-g_{1})\beta(x,z)$ in equation (50) may help us to differentiate two DGPs by generating a differently-signed HTE.
@@
If, under certain $x$ and $z$, (51) and the first term of (50) have the same sign and the second term of (50) has the opposite sign and is sufficiently large, (51) and (50) will have different signs. Moreover, If $e$ is uniformly distributed, then $f^{\prime}_{e}=0$ and thus equation (51) is equal to 0 while equation (50) is non-zero.
@@
Consider the transformed outcome $h(Y_{s})$ satisfying equation (47) in which moderation effect $\beta(x,z)$ is monotonic.
@@
Proposition 12 reveals and Corollary A1 illustrates that Assumptions 1-2 and monotonicity are not, in general, jointly sufficient to use HTEs to assess mechanism activation on tansformed outcomes.
H.2 Assumptions on the Latent Utility Distributions
In practice, one of the most common cases is that $Y$ is the latent utility, and $h(Y)$ is the observed action or discrete choice. If we can re-construct the utility from the observed data, then we can use it as the new outcome and use HTEs to assess mechanism activation under Assumptions 1-2. From Propositions 2 and 3, we know more information on mechanism activation can be ascertained from HTEs on the latent outcome.
@@ H.2 Assumptions on the Latent Utility Distributions
In practice, utility is usually specified as a linear function $V^{i}_{m}=X\beta$, where $X$ is the vector of observed variables and $\beta$, are parameters. We can estimate $\beta$ from the data. Then, we could treat $V_{m}^{i}$ as the new outcome to explore the mechanism using HTE. It should be clear that Propositions 2 and 3 hold for $V_{i}^{m}$ even if the utility is $U$ by linearity.
Now, we use our motivating example to illustrate how to apply the above RUM and distribution assumptions in practice. In the main text, we provide a simple model of the effect of corruption information on pro-incumbent voting. The model specifies the systematic component of utility, $V_{m}^{i}$. A RUM requires us to specify a random component of utility, $\varepsilon$ which also affects individual behavior and is additively separable from $V_{m}^{i}$ (i.e. not relevant to the theory of interest). Therefore, given the observed voting data we have, if we assume the distribution of $\varepsilon$ is type-I extreme value, then the probability individual votes for the incumbent is
@@
In practice, researchers tend to specify a linear model to approximate $V_{m}^{i}$, though other functional forms are also possible. If $V_{m}^{i}$ is correctly specified, we then treat $V_{m}^{i}$ as an outcome (utility) and use this measure when estimating HTE.
H.3 Assumptions about the Magnitude of HTEs
@@
$\mathbb{P}[M_{1}\text{ is active}\mid\text{as extreme as HTE}]$ $=\frac{p\mathbb{P}[\text{HTE}|\text{active}]}{p\mathbb{P}[\text{HTE}|\text{active}]+(1-p)\mathbb{P}[\text{HTE}|\text{not active}]}$
@@
[…]